# Elevating Software Trust: Unveiling and Quantifying the Risk Landscape

### v1

@@ Abstract
Considering the ever-evolving threat landscape and rapid changes in software development, we propose a risk assessment framework SRiQT (Software Risk Quantification through Trust). This framework is based on the necessity of a dynamic, data-driven, and adaptable process to quantify risk in the software supply chain. Usually, when formulating such frameworks, static pre-defined weights are assigned to reflect the impact of each contributing parameter while aggregating these individual parameters to compute resulting risk scores. This leads to inflexibility, a lack of adaptability, and reduced accuracy, making them unsuitable for the changing nature of the digital world. We adopt a novel perspective by examining risk through the lens of trust and incorporating the human aspect. Moreover, we quantify risk associated with individual software by assessing and formulating risk elements quantitatively and exploring dynamic data-driven weight assignment. This enhances the sensitivity of the framework to cater to the evolving risk factors associated with software development and the different actors involved in the entire process. The devised framework is tested through a dataset containing 9000 samples, comprehensive scenarios, assessments, and expert opinions. Furthermore, a comparison between scores computed by the OpenSSF scorecard, OWASP risk calculator, and the proposed SRiQT framework has also been presented. The results suggest that SRiQT mitigates subjectivity and yields dynamic data-driven weights as well as risk scores.
@@
Risk Quantification, Risk in Software Supply Chain, Software Risk Score, Risk Assessment Framework, Trust Perspective on Software Risk.
@@ I Introduction
The diversity of entities engaged in the process, along with the 3rd party components (e.g., equipment, code, and libraries), magnifies the susceptibilities and risks [4, 5]. According to the European Union Agency for Cybersecurity’s (ENISA) report, attacks related to the software supply chain are considered the prime threat in the current threat landscape [6]. Vulnerability in or attack on even a single component of these chains has catastrophic impacts as it has the potential to propagate to other components within the same supply chain [7]. Therefore, it is of paramount importance that such risks be identified and addressed suitably. Moreover, if there exists multiple vulnerabilities or risks, an organization has to determine which vulnerability to address first [8].
@@
When choosing software to download, users must compare multiple (risky) options and make decisions based on objective criteria rather than subjective opinion [9]. Risk assessment guides decisions in scenarios like prioritizing resource allocation for addressing security issues or choosing safer software options [8]. Moreover, it is important to quantify the risk to estimate the severity of the risk and eliminate subjectivity and inconsistencies in evaluating these risk levels to make informed decisions [10]. For instance, a subjective risk assessment performed to evaluate a piece of software might yield a low risk level when carried out by an individual, and the same piece of software might be assessed as a high risk level by another individual in the same setting in the same organization. This subjective and inconsistent assessment makes it difficult for the organization to make decisions regarding the software.
@@
Figure 1: Proposed SRiQT Framework: Individual parameters are computing the actor-based risks (i.e., developer, publisher, and the user) and their respective weights, which are fed along with the penalty factor to compute the final risk score.
@@
Most research works on software risk assessment centers around qualitative analyses [14, 15, 16, 17, 18, 19], which introduces a factor of subjectivity and inconsistency in outcomes. For instance, if two people evaluating the risk associated with the same software have different levels of experience/expertise, the risk assessed by them will be different due to their subjective opinions. In the context of quantitative risk evaluation, weights are associated with individual parameters to reflect their influence on the cumulative risk score. These weights are often static, pre-defined, and there exists a lack of justification for the selection of preset values for these weights (e.g., if the number of parameters influencing the cumulative risk score is $p$, taking a mean of all these parametric values suggest a static and pre-defined weight of $\frac{1}{p}$ associated with each parameter without any reasoning to justify why the equal weight $\frac{1}{p}$ is employed instead of $\frac{0.5}{p}$ for one and $\frac{1.5}{p}$ for the other and so on), which also introduces an element of subjectivity. The objective quantification of these weights has been overlooked while formulating risk evaluation models [14, 20]. Moreover, a significant proportion of these risk assessment frameworks emphasize only the software development stage and accordingly employ parameters regarding only the software code itself [21, 22]. In addition, the omission of software dependencies from risk management models is a significant limitation that potentially undermines the effectiveness of these models [23]. Furthermore, the human aspect has been ignored while defining the threat landscape for risk assessment.
@@
Trust and risk are like two peas in a pod; assessing the trustworthiness of an entity helps in understanding the potential risks associated with that entity [11]. In other words, trust is seen as a risk mitigation strategy, and it reduces perceived risk [12]. The existing literature indicates a noteworthy gap in examining software risk, specifically in relation to the perspective of trust. To the best of our knowledge, none of the current studies have explored software risk from the trust standpoint, highlighting a substantial and unexplored area within the research landscape. It is of significant importance to study and define trust to assess software ecosystems [13]. How trust dynamics interplay with software-related risks is worth investigating and will provide valuable insights into the vulnerabilities and challenges faced in the domain of software supply chain.
@@
Accordingly, we propose a novel software risk assessment framework SRiQT (Software Risk Quantification through Trust) that relies on quantitative influencing parameters to formulate individual risk segments along with data driven and dynamic weights associated with these risk segments to address subjectivity and inconsistency in risk assessment. SRiQT focuses on the human aspect of the software supply chain and incorporates parameters related to the developer, publisher, and the user in addition to software code and dependencies-related attributes from a trust perspective, i.e., it assesses the risk associated with the software through the parameters acquired from different parts of the software supply chain that are based on the major components from the trust management domain such as reputation, experience, recommendation, penalty, etc. In other words, the computation of risks associated with software and the subsequent decision-making processes are approached and analyzed through the lens of trust. Furthermore, it accounts for context, integrates historical occurrences, and leverages timeliness to manage risk propagation. Figure 1 outlines the proposed SRiQT framework.
I-A Our Contributions
The research at hand focuses on the following main contributions:
@@ I-A Our Contributions
  • Envisaging a software risk assessment framework from the trust viewpoint SRiQT (Software Risk Quantification through Trust), which has not been undertaken previously to the best of our knowledge.
@@
  • Introducing novel non-subjective attributes and quantifying these attributes and individual risk segments to yield consistent and non-subjective risk scores which the existing tools such as the Open Web Application Security Project (OWASP) risk calculator[25] lack. Please note that subjectivity here does not refer to context-based subjectivity but bias or partiality.
  • Employing suitable and rational influencing parameters as weights and catering to up-to-date/current risk factors while formulating contributing parameters to address the problem of weight quantification and achieving data driven and dynamic weights associated with individual attributes and risk segments. This feature is lacking in the existing tools such as the Open Source Security Foundation (OpenSSF) scorecard [26] and OWASP risk calculator.
@@
I-B Organization of the Paper
The rest of the paper is organized as follows. Section II reviews the state-of-the-art risk assessment models and tools. Section III presents the problem statement, provides definitions for the major components, and specifies participating actors, Section IV examines the step-wise rationale behind the selected influencing parameters, Section V covers the details of our proposed SRiQT’s system model, whereas Section VI discuses the parameter tuning for tackling particular scenarios and tailoring specific elements within the envisaged SRiQT framework. Section VII offers insights into the dataset that has been employed and reports the results, and Section VIII presents a discussion regarding the claimed contributions, yielded results, and comparisons. Finally, Section IX concludes the paper and outlines future directions.
@@ I-B Organization of the Paper
OWASP
OpenSSF Scorecard
CVSS
Risk Explorer
EU-Risk
SRiQT
@@ II Related Work
This section presents a detailed literature review of the existing risk assessment works in both academia and the industry, followed by an overview of the contrasting features among the current tools (i.e., OpenSSF scorecard and OWASP) and the proposed SRiQT framework. Table I exhibits a comparison between existing risk assessment tools in the industry and the proposed SRiQT framework.
@@ II-A Literature Review
The OWASP risk rating methodology utilizes a conventional risk assessment approach, wherein risk is determined by both the probability of an attack occurring and the potential impact of a successful attack [25]. However, it introduces subjectivity, employs pre-defined scores to qualitative factors without a data-driven approach which is difficult to follow (e.g., score mappings of 1, 4, and 9 for motives), the overall likelihood and impact scores have been computed by averaging out all the scores for individual factors which implies that all the considered factors influence the final risk score equally.
@@
Table II: Comparison - SRiQT with OWASP Risk Calculator [36] and OpenSSF Scorecard [35].
Features OWASP OpenSSF SRiQT
@@ II-B Existing Tools
We compare the proposed SRiQT framework with the existing security/risk calculation tools OpenSSF scorecard and OWASP risk calculator by highlighting the differences and further explaining the need for the envisaged SRiQT framework. Refer to Table II for details.
@@ III Problem setup
Risk is defined as the distrust of an entity towards a target entity or an entire system relying on context, reputation, and recommendations/opinions.
@@ III-1 Developer
The developer, playing a central role among the key participants, is crucial for shaping the software through involvement in its design, coding, and comprehensive development. The quality, functionality, and security of the software product are directly impacted by the developer’s expertise and decision-making throughout the development phase.
@@ IV Parameter Selection
Selecting the appropriate parameters for risk assessment of software from a trust perspective is a critical step in devising a trust-driven risk evaluation framework for ensuring the security of software systems. It is imperative to consider a multitude of factors that impact trust and security when determining the parameters that contribute to this assessment. This encompasses a blend of standards’ requirements, published peer-reviewed work, expert consultation, and observable attributes associated with the software.
@@ V System Model
Let an individual software be $S_{i}$, wherein $i\in\{1,2,3,...,I\}$, and a time instance is denoted by $t_{k}$, where $k\in\{1,2,3,...,T\}$. Please note that the time index is being used to acknowledge the dynamic nature of risks, which can change with evolving circumstances or influencing risk factors. It is crucial to know when the risk computation took place, or in other words, how old the risk assessment is.
@@ V-A Developer based Risk:
The developer-based risk ($R^{\textup{DEV}}_{S_{i},t_{k}}$) is an amalgamation of the risk from code dependencies, code specifications, programming language, historical information on vulnerabilities, and the developer’s experience associated with the software $S_{i}$ at time instance $t_{k}$.
@@ V-A1 Code Dependencies
Motivation: Literature and security solution providers indicate that a higher number of dependencies increases risk and suggest minimizing dependencies to mitigate associated risks [50, 51]. It is worth noting that to minimize subjectivity, factors such as the quality of dependencies have been omitted from consideration.
@@ V-A2 Code Specifications
Motivation: Literature suggests the reliance of software risk on the code characteristics [22, 52]. Moreover, studies indicate the impact of the developer’s reputation on different phases of the software review process [53]. Multiplication operation is employed to capture the proportion contribution of the code characteristics and to intensify the effect of vulnerabilities in $R^{\textup{CS}}_{S_{i},t_{k}}$.
@@
$w^{\textup{LC}}_{t_{k}}=\frac{\sum_{i=1}^{I}V^{\textup{D}}_{S_{i},t_{k}}}{|S^{\textup{D}}_{t_{k}}|},$ (3)
where $V^{\textup{D}}_{S_{i},t_{k}}$ is the number of vulnerabilities in all the software developed by the same developer, and $S^{\textup{D}}_{t_{k}}$ is the set of all software developed by the same developer at time $t_{k}$. Simply put, the weight $w^{\textup{LC}}_{t_{k}}$ represents the average number of vulnerabilities per software developed by a developer. It is worth noting that for the developer who has previously developed software with no vulnerabilities, the weight $w^{\textup{LC}}_{t_{k}}$ will be zero regardless of the total number of software developed. In other words, even if the developer has developed only a couple of software before and no vulnerability has been reported on that software, the developer will be given a chance with the code specification-based risk.

### v2

@@ Abstract
Considering the ever-evolving threat landscape and rapid changes in software development, we propose a risk assessment framework called SAFER (Software Analysis Framework for Evaluating Risk). This framework is based on the necessity of a dynamic, data-driven, and adaptable process to quantify security risk in the software supply chain. Usually, when formulating such frameworks, static pre-defined weights are assigned to reflect the impact of each contributing parameter while aggregating these individual parameters to compute resulting security risk scores. This leads to inflexibility, a lack of adaptability, and reduced accuracy, making them unsuitable for the changing nature of the digital world. We adopt a novel perspective by examining security risk through the lens of trust and incorporating the human aspect. Moreover, we quantify security risk associated with individual software by assessing and formulating risk elements quantitatively and exploring dynamic data-driven weight assignment. This enhances the sensitivity of the framework to cater to the evolving security risk factors associated with software development and the different actors involved in the entire process. The devised framework is tested through a dataset containing 9000 samples, comprehensive scenarios, assessments, and expert opinions. Furthermore, a comparison between scores computed by the OpenSSF scorecard, OWASP risk calculator, and the proposed SAFER framework has also been presented. The results suggest that SAFER mitigates subjectivity and yields dynamic data-driven weights as well as security risk scores.
@@
Risk Quantification, Risk in Software Supply Chain, Software Security Risk Score, Risk Assessment Framework, Trust Perspective on Software Risk.
@@ I Introduction
The diversity of entities engaged in the process, along with the 3rd party components (e.g., equipment, code, and libraries), magnifies the susceptibilities and security risks [4, 5]. According to the European Union Agency for Cybersecurity’s (ENISA) report, attacks related to the software supply chain are considered the prime threat in the current threat landscape [6]. Vulnerability in or attack on even a single component of these chains has catastrophic impacts as it has the potential to propagate to other components within the same supply chain [7]. Therefore, it is of paramount importance that such security risks be identified and addressed suitably. Moreover, if there exist multiple vulnerabilities or security risks, an organization has to determine which vulnerability to address first [8]. Please note that the term risk refers to security risk throughout the paper.
@@
When choosing software to download, users must compare multiple (risky) options and make decisions based on objective criteria rather than subjective opinion [9]. Risk assessment guides decisions in scenarios like prioritizing resource allocation for addressing security issues or choosing safer software options [8]. Moreover, it is important to quantify the risk to estimate the severity of the risk and eliminate subjectivity and inconsistencies in evaluating these risk levels to make informed decisions [10]. For instance, a subjective risk assessment performed to evaluate a piece of software might yield a low-risk level when carried out by an individual, and the same piece of software might be assessed as a high-risk level by another individual in the same setting in the same organization. This subjective and inconsistent assessment makes it difficult for the organization to make decisions regarding the software.
@@
Figure 1: Proposed SAFER Framework: Individual parameters are computing the actor-based risks (i.e., developer, publisher, and the user) and their respective weights, which are fed along with the penalty factor to compute the final risk score.
@@
Most research works on software risk assessment centers around qualitative analyses [14, 15, 16, 17, 18, 19], which introduces a factor of subjectivity and inconsistency in outcomes. For instance, if two people evaluating the risk associated with the same software have different levels of experience/expertise, the risk assessed by them will be different due to their subjective opinions. In terms of quantitative risk evaluation, weights are associated with individual parameters to reflect their influence on the cumulative risk score. These weights are often static, pre-defined, and there exists a lack of justification for the selection of preset values for these weights (e.g., if the number of parameters influencing the cumulative risk score is $p$, taking a mean of all these parametric values suggest a static and pre-defined weight of $\frac{1}{p}$ associated with each parameter without any reasoning to justify why the equal weight $\frac{1}{p}$ is employed instead of $\frac{0.5}{p}$ for one and $\frac{1.5}{p}$ for the other and so on), which also introduces an element of subjectivity. The objective quantification of these weights has been overlooked while formulating risk evaluation models [14, 20]. Moreover, a significant proportion of these risk assessment frameworks emphasize only the software development stage and accordingly employ parameters regarding only the software code itself [21, 22]. In addition, the omission of software dependencies from risk management models is a significant limitation that potentially undermines the effectiveness of these models [23]. Furthermore, the human aspect has been ignored while defining the threat landscape for risk assessment.
@@
The perceived reliability and security of software considerably impact the willingness of an end user to use it, i.e., if the software is susceptible to security risks, the user’s confidence in that software weakens. Assessing the trustworthiness of an entity helps in understanding the potential risks associated with that entity [11]. In other words, trust is seen as a risk mitigation strategy, and it reduces perceived risk [12]. The existing literature indicates a noteworthy gap in examining software risk, specifically in relation to the perspective of trust. To the best of our knowledge, none of the current studies have explored software risk from the trust standpoint, highlighting a substantial and unexplored area within the research landscape. It is of significant importance to study and define trust to assess software ecosystems [13].
@@
Accordingly, we propose a novel software risk assessment framework called SAFER (Software Analysis Framework for Evaluating Risk) that relies on quantitative influencing parameters to formulate individual risk segments along with data-driven and dynamic weights associated with these risk segments to address subjectivity and inconsistency in risk assessment. It is essential to emphasize that this paper considers risk from a security point of view. While referring to dynamic and data-driven weights, it is worth mentioning that the weights themselves do not impact secure software development practices but reflect/represent the characteristics of the software development and vulnerabilities. Employing data-driven weights adds to the transparency of the risk assessment by reflecting the actual threat landscape. Moreover, if an end user (with limited or no knowledge regarding software development or exposure) wishes to customize the framework, they will require a deep understanding of each and every aspect to make an informed decision regarding the selection of appropriate pre-defined static weights. Whereas, in the proposed framework, these weights are kept dynamic so they will automatically adapt to the data provided without needing a thorough understanding. SAFER also considers the human aspect of the software supply chain and incorporates parameters related to the developer, publisher, and the user in addition to software code and dependencies-related attributes from a trust perspective, i.e., it assesses the risk associated with the software through the parameters acquired from different parts of the software supply chain that are based on the major components from the trust management domain such as reputation, experience, recommendation, penalty, etc. In other words, the computation of risks associated with software and the subsequent decision-making processes are approached and analyzed through the lens of trust. Furthermore, it accounts for context, integrates historical occurrences, and leverages timeliness to manage risk propagation. Figure 1 outlines the proposed SAFER framework. The proposed framework is expected to have practical implications, including but not limited to (i) informed decision making, organizations and individual users can make well-informed choices about which software to adopt, (ii) reducing security incidents by assisting users in identifying low risk software, (iii) risk reduction for organizations by identifying potential risk, and (iv) reassuring adoption of best security practices by highlighting the susceptible areas of the software supply chain.
Accordingly, the research at hand focuses on the following main contributions:
@@
  • Envisaging a software risk assessment framework from the trust viewpoint, SAFER (Software Analysis Framework for Evaluating Risk), which has not been undertaken previously to the best of our knowledge.
@@
  • Introducing novel non-subjective attributes (i.e., attributes that can be measured or assessed without personal bias or opinions) and quantifying these attributes and individual risk segments to yield consistent and non-subjective risk scores which the existing tools such as the Open Web Application Security Project (OWASP) risk calculator[25] lack. Please note that subjectivity here does not refer to context-based subjectivity but bias or partiality.
  • Employing suitable and rational influencing parameters as weights and catering to up-to-date/current risk factors while formulating contributing parameters to address the problem of weight quantification and achieving data-driven and dynamic weights associated with individual attributes and risk segments. This feature is not available in the existing tools such as the Open Source Security Foundation (OpenSSF) scorecard [26] and OWASP risk calculator.
@@
The rest of the paper is organized as follows. Section II reviews the state-of-the-art risk assessment models and tools. Section III presents the problem statement, provides definitions for the major components, and specifies participating actors, Section IV examines the step-wise rationale behind the selected influencing parameters, Section V covers the details of our proposed SAFER’s system model, whereas Section VI discuses the parameter tuning for tackling particular scenarios and tailoring specific elements within the envisaged SAFER framework. Section VII offers insights into the dataset that has been employed and reports the results, and Section VIII presents a discussion regarding the claimed contributions, yielded results, and comparisons. Finally, Section IX concludes the paper and outlines future directions.
@@
OWASP [25]
OpenSSF Scorecard [26]
CVSS [14]
Risk Explorer [29]
EU-Risk [28]
SAFER
@@ II Related Work
This section presents a detailed literature review of the existing risk assessment works in both academia and the industry, followed by an overview of the contrasting features among the current tools (i.e., OpenSSF scorecard, and OWASP) and the proposed SAFER framework. Table I exhibits a comparison between existing risk assessment tools in the industry and the proposed SAFER framework. The calculators and risk evaluation methods mentioned were employed to assess risk, and the underlying literature on their design/development and functionality was thoroughly studied to deduce the conclusions presented in Table I.
@@ II-A Literature Review
The OWASP risk rating methodology utilizes a conventional risk assessment approach, wherein risk is determined by both the probability of an attack occurring and the potential impact of a successful attack [25]. However, it introduces subjectivity, employs pre-defined scores to qualitative factors without a data-driven approach, which is difficult to follow (e.g., score mappings of 1, 4, and 9 for motives), the overall likelihood and impact scores have been computed by averaging out all the scores for individual factors which implies that all the considered factors influence the final risk score equally.
@@
Table II: Comparison - SAFER with OWASP Risk Calculator [36] and OpenSSF Scorecard [35].
Features OWASP OpenSSF SAFER
@@ II-B Existing Tools
We compare the proposed SAFER framework with the existing security/risk calculation tools OpenSSF scorecard and OWASP risk calculator by highlighting the differences and further explaining the need for the envisaged SAFER framework. Refer to Table II for details.
@@ III Problem setup
According to ISO 31000, risk is defined as the impact of uncertainty on achieving objectives. Here, the objectives refer to software security. It can be described as a combination of the potential consequences of an event and the likelihood of it occurring [41]. The higher the risk associated with a target entity, the higher the distrust of an entity towards that target entity becomes.
@@ III-1 Developer
The developer, playing a central role among the key participants, is crucial for shaping the software through involvement in its design, coding, and comprehensive development. The developer’s expertise and decision-making directly impact the quality, functionality, and security of the software product throughout the development phase.
@@ IV Parameter Selection
Selecting the appropriate parameters for risk assessment of software from a trust perspective is a critical step in devising a trust-driven risk evaluation framework for ensuring the security of software systems. It is imperative to consider a multitude of factors that impact trust and security when determining the parameters that contribute to this assessment. For instance, (i) the parameters should be quantifiable, which means that each parameter should easily be expressed in numeric values; (ii) the parameters must be non-subject, which means that their numeric value can be measured objectively without any personal bias or opinions; (iii) the statistics of the software should be available (publicly available for the evaluation of open source software) which means that the values for each parameter must be accessible to the evaluator; (iv) the parameters should be simple for user-friendly calculation process, and (v) parameters like cost of software are not considered as cost does not impact the suitability of a piece of software as a less or more risky option. The process of cataloging various impacting parameters encompasses a blend of standards’ requirements, published peer-reviewed work, expert consultation, and observable attributes associated with the software. Moreover, a structured evaluation of suitability attributes has been performed by employing the logic scoring of preferences (LSP) approach [51].
@@ IV-D Step 4: Observable Attributes
IV-E Step 5: Logic Scoring of Preferences
Logic scoring of preferences (LSP) has been employed as a structured approach to evaluate suitability attributes. The primary objective of such methods is to offer a quantitative approach for making evaluation decisions based on criteria that accurately represent stakeholder goals and requirements. LSP works on the fundamental notion of completely aligning with human intuitive processes of evaluation and selection [51]. The suitability attributes for assessing the software security in the proposed framework do not include any cost-related attributes, as the cost does not impact the suitability of a piece of software as a less or more risky option. The stepwise decomposition method to develop the suitability attribute tree starts with the identification of four categories of attributes, including (i) reputation/experience, which contains parameters related to the experience and reputation of the actors (i.e., developer’s experience in programming language, the developer’s expertise, and the publisher’s experience) (ii) exposure, which includes parameters related to susceptibility (i.e., dependencies, code coverage, unresolved vulnerabilities, and resolved vulnerabilities), (iii) popularity and maintenance, which covers parameters that describe the ongoing development and acceptance of software (i.e., frequency updates, rating, and downloads), and (iv) supplements, which encompasses attributes whose changing value will not enhance nor weaken the security of software, however, these can help in the assessment of the security of a software in conjunction of other parameters (i.e., code length, forks, and context).
@@ V System Model
Let an individual software be $S_{i}$, wherein $i\in\{1,2,3,...,I\}$, and a time instance is denoted by $t_{k}$, where $k\in\{1,2,3,...,T\}$. Please note that the time index is being used to acknowledge the dynamic nature of risks, which can change with evolving circumstances or influencing risk factors. It is crucial to know when the risk computation took place, or in other words, how old the risk assessment is. As mentioned in Section IV, the proposed risk assessment framework employs LSP for the suitability attributes’ aggregation. The overall risk score has been divided into three risk segments corresponding to each actor. Within these risk segments, the LSP concept of maximum substitutability or orness has been utilized, and all the degrees of orness/andness have been decided by data-driven weights. Moreover, all suitability attributes are objectively quantifiable, independent, and can be measured individually. In addition, it is noteworthy that the notion of substitutability also brings in the property of compensativeness, which helps accommodate the unavailability of data against any parameter(s). Furthermore, if a risk segment cannot be evaluated due to limited or no data availability, or due to an undefined output, it must be assumed that this particular risk segment carries the maximum possible risk (i.e., normalized value equal to 1). In case there is data available only for some of the parameters within a risk segment data, the remaining parameters can be put as zero while computing that risk segment.
@@ V-A Developer based Risk:
The developer-based risk ($R^{\textup{DEV}}_{S_{i},t_{k}}$) is an amalgamation of the risk from code dependencies, code specifications, programming language, historical information on vulnerabilities, and the developer’s experience associated with the software $S_{i}$ at time instance $t_{k}$. The LSP property of compensativeness applies to all three risk segments that fall under the developer-based risk.
@@ V-A1 Code Dependencies
Motivation: Literature and security solution providers indicate that a higher number of dependencies increases risk and suggest minimizing dependencies to mitigate associated risks [52, 53]. It is worth noting that to reduce subjectivity, factors such as the quality of dependencies have been omitted from consideration as suitability attributes while applying the LSP process.
@@ V-A2 Code Specifications
Motivation: Literature suggests the reliance of software risk on the code characteristics [22, 54]. Moreover, studies indicate the impact of the developer’s reputation on different phases of the software review process [55]. Multiplication operation is employed to capture the proportion contribution of the code characteristics and to intensify the effect of vulnerabilities in $R^{\textup{CS}}_{S_{i},t_{k}}$. As the code length belongs to the supplement category of suitability attributes in LSP, the property of simultaneity has been applied here.
@@
$w^{\textup{LC}}_{t_{k}}=\frac{\sum_{j}\sum_{i=1}^{I}V^{\textup{D}}_{j,S_{i},t_{k}}}{\sum_{j}|S^{\textup{D}}_{j,t_{k}}|},$ (3)
where $V^{\textup{D}}_{S_{i},t_{k}}$ is the number of vulnerabilities in all the software developed by the same developer, $j$ denotes a developer, and $S^{\textup{D}}_{t_{k}}$ is the set of all software developed by the same developer at time $t_{k}$. Simply put, the weight $w^{\textup{LC}}_{t_{k}}$ represents the average number of vulnerabilities per software developed by a developer. It is worth noting that for the developer who has previously developed software with no vulnerabilities, the weight $w^{\textup{LC}}_{t_{k}}$ will be zero regardless of the total number of software developed. In other words, even if the developer has developed only a couple of software before and no vulnerability has been reported on that software, the developer will be given a chance with the code specification-based risk.